The United States Department of Agriculture reports that, under its standard inspection system one in every 100 slaughtered chickens passes inspection for faecal contamination. Find the probability that, in a random sample of 6 slaughtered chickens, at least one passes inspection for faecal inspection.


1
Expert's answer
2022-06-12T14:54:33-0400

Let "X=" the number slaughtered chickens passed inspection for faecal contamination: "X\sim Bin (n, p)."

Given "n=6, p=0.01., q=1-p=0.99."


"P(X\ge 1)=1-P(X=0)"

"=1-\dbinom{6}{0}(0.01)^{0}(0.99)^{6-0}"

"=0.058519850599"